OpenStudy (kitkat1):

A set of equations is given below: Equation S: y = x + 9 Equation T: y = 2x + 1 Which of the following steps can be used to find the solution to the set of equations? x + 9 = 2x + 1 x = 2x + 1 x + 9 = 2x x + 1 = 2x + 9

OpenStudy (kitkat1):

Medals!

OpenStudy (kitkat1):

Help please will give medals

OpenStudy (anonymous):

u go to conections 8grade

OpenStudy (kitkat1):

the quiz is 08.03

OpenStudy (anonymous):

so u go to connections 8grade

OpenStudy (kitkat1):

yes 8th grade

OpenStudy (anonymous):

connections academy?

OpenStudy (kitkat1):

yes

OpenStudy (anonymous):

msg me>

OpenStudy (kitkat1):

do u know the answer?

OpenStudy (anonymous):

wait if you go to the same school msg me do u take pre algebra?

OpenStudy (kitkat1):

yes pre algebra, why cant u tell me the answer?

OpenStudy (anonymous):

msg me then i wil tell?

OpenStudy (kitkat1):

how can I do that?

OpenStudy (anonymous):

on the top the envelope is the messaging icon click on it read the msg then green arrow to reply

OpenStudy (misty1212):

HI!!

OpenStudy (kitkat1):

Will u be able to help me?

OpenStudy (misty1212):

set them equal, get \[x+9=2x+1\]

OpenStudy (kitkat1):

okay can u help with one more similar one?

OpenStudy (kitkat1):

A system of equations is given below: y = 3x + 4 2x + 3y = 10 Which of the following steps could be used to solve by substitution? (4 points) 2(3x + 4) + 3y = 10 3(y = 3x + 4) 2x + 3(3x + 4) = 10 3x + 4 = 2x + 3y

OpenStudy (kitkat1):

Mitsy1212

OpenStudy (kitkat1):

Misty1212